### v1

@@ Appendix A Omitted proofs
This appendix contains the proofs that were deferred from the main text. Section A.1 gives algebraic and embedding lemmas that support the analysis of the cut-matching game. In Section A.2 we introduce the fair-cut framework of [7] and use it to analyze flow-based routines: first the flow subroutine in the cut-matching game (Lemma 3.6), and then the trimming step (Theorem 5.1).
@@ A.1 Algebraic and Embedding Proofs
Case $P_{L}\geq\frac{1}{20}P_{A_{t}}$: We set $\eta=0$, $A_{t}^{r}=R$, and $\bar{m}_{i}=\mu(i)$ for $i\in A_{t}^{r}$. The construction of $A_{t}^{l}$ is as follows: If $\mu(L)\leq\mu(A_{t})/8$, then we set $A_{t}^{l}=L$, and $m_{i}=\mu(i)$ for $i\in A_{t}^{l}$.
Otherwise, we sort $L$ according to $u_{i}$ and scan the vertices (from smallest to largest) until the accumulative $\mu(\cdot)$ weight of the scanned vertices is at least $\mu(A_{t})/8$. Let $i_{1},\ldots,i_{k}$ be the scanned vertices. We set $A_{t}^{l}=\{i_{1},\ldots,i_{k}\}$, $m_{i_{j}}=\mu(i_{j})$ for $1\leq i_{j}<k$, and $m_{i_{k}}=\mu(A_{t})/8-\mu(A_{t}^{l}\setminus\{i_{k}\})$. It is easy to verify that properties (1)(1)-$(3)$ of the lemma hold. Property $(4)$ holds because of the choice $\eta=0$. Finally, property $(5)$ holds since $P_{L}\geq\frac{1}{20}P_{A_{t}}$ and since $A_{t}^{l}$ contains the vertices with the smallest $u_{i}$ values (of total $\mu(\cdot)$ weight equaling $\mu(A_{t})/8$, while $\mu(L)\leq\mu(A_{t})/2$) in $L$.
@@
  • $A_{t}^{\ell}$ is the subset of $R^{\prime}$ consisting of the vertices with the largest $u_{i}$ values, chosen so that their total $\mu(\cdot)$ weight satisfies $\mu(A_{t}^{l})=\mu(A_{t})/8$. [^14] This is done similarly to the previous case. Again, $m_{i}=\mu(i)$ for every $i\in A_{t}^{l}$, except for possibly one vertex.
@@ A.2 Fair Cuts
In this section, we prove Lemma 3.6 and Theorem 5.1, concerning the flow routines in the cut-matching step and the trimming step. The proofs are based on fair cuts, a stronger notion of approximate min-cut, introduced by [7].
Definition A.1 (Fair Cut, [7, 6]).
@@
Theorem A.2 (Fair Cut, Theorem 1.1 of [7]).
@@
Let $\alpha=0.1$. Consider the following $(s,t)$-flow problem on a new graph $H=(V_{H}=V\cup\{s,t\},E_{H},w_{H})$ constructed as follows. We begin with $G$ and set its edge capacities to $\frac{c}{1+\alpha}$. We then add the source $s$ and connect it to every $v\in A^{l}$ with capacity $m_{v}$. Finally, we add the sink $t$ and connect it to every $v\in A^{r}$ with capacity $\frac{1}{1+\alpha}\bar{m}_{v}$.
@@
We compute a $(1+\alpha)$-fair cut $(S,T)$ with a corresponding feasible flow $f$ in $\tilde{O}(m)$ time. We split into the following two cases:
@@
Case 1: $S=\{s\}.$ By the definition of fair cut, for every $v\in A^{l}$, $f$ sends from $s$ to $v$ at least $\frac{1}{1+\alpha}m_{v}$ flow. By decomposing $f$ into flow paths and reducing the flow sent on each path, we obtain a flow $f^{\prime}$ for $H$ that sends from $s$ to $v$ exactly $\frac{1}{1+\alpha}m_{v}$ flow, for every $v\in A^{l}$. We claim that $f^{\prime\prime}=(1+\alpha)f^{\prime}$ is a feasible flow for $\Pi(G)$. Indeed, in $f^{\prime\prime}$, every $v\in A^{l}$ receives $m_{v}=$ flow from $s$, every $v\in A^{r}$ sends at most $(1+\alpha)\cdot\frac{1}{1+\alpha}\bar{m}_{v}=\bar{m}_{v}$ flow to $t$, and every $e\in E$ carries at most $(1+\alpha)\cdot\frac{1}{1+\alpha}c=c$ units of flow.
@@
Case 2: $S\neq\{s\}.$ We split the cut edges $E_{H}(S,T)$ into three sets: (i) the edges of $G$ crossing the cut $E_{m}=E_{G}(S\setminus\{s\},T\setminus\{t\})$, (ii) the cut edges incident to $s$ $E_{s}=\{(s,v)\in E_{H}\mid v\in T\}$, and (iii) the cut edges incident to $t$ $E_{t}=\{(v,t)\in E_{H}\mid v\in S\}$. Let $\bar{E}_{s}$ be the incident edges to $s$ that are not in $E_{s}$.
@@
We decompose $f$ into paths. By the definition of a fair cut, every edge $(s,v)\in E_{s}$ sends at least $\frac{1}{1+\alpha}m_{v}$ flow. We scale down the flow in the flow paths as follows. We delete all flow paths that start with an edge of $\bar{E}_{s}$. The rest of the flow paths are scaled down until each edge $(s,v)\in E_{s}$ sends exactly $\frac{1}{(1+\alpha)}m_{v}$ flow. This results in a flow $f^{\prime}$. The algorithm then returns $f^{\prime\prime}=(1+\alpha)f^{\prime}$ and the cut $S^{\prime}=S\setminus\{s\}$.
@@
Similarly to the previous case, $f^{\prime\prime}$ is a feasible flow for the flow problem $\Pi(G\setminus S^{\prime})$. It remains to prove that $\Phi_{G}^{\mu}(S^{\prime},V\setminus S^{\prime})\mathrel{\mathop{:}}=\frac{|E_{m}|}{\min\{\mu(S^{\prime}),\mu(V\setminus S^{\prime})\}}\leq\frac{7}{c}$. We begin by upper bounding the numerator. For the remainder of the analysis, we consider again the flow $f$ and its path decomposition. By the definition of a fair cut, every flow path that traverses through $E_{m}$ must start with an edge in $\bar{E}_{s}$, and every edge in $E_{m}$ is at least $\frac{1}{1+\alpha}$ saturated. Therefore
$\frac{c}{1+\alpha}\cdot|E_{G}(S^{\prime},V\setminus S^{\prime})|=w_{H}(E_{m})\leq(1+\alpha)w_{H}(\bar{E}_{s})=(1+\alpha)m(S^{\prime})\leq(1+\alpha)\mu(S^{\prime}).$ (5)
@@
To bound the denominator we prove that $\mu(V\setminus S^{\prime})=\Omega(\mu(S^{\prime}))$. Recall the assumption of the lemma $\bar{m}(A^{r})\geq\mu(V)/2$. We prove that $\bar{m}(A^{r}\cap S^{\prime})$ is small. Indeed, for every $v\in A^{r}\cap S^{\prime}$, $(v,t)\in E_{t}$ is crossing the cut $(S,T)$ and therefore is $\frac{1}{1+\alpha}$ saturated. Since the total flow crossing $(S,T)$ is bounded by the total source capacity (which is $m(A^{l})$), we get that
$\frac{1}{1+\alpha}\bar{m}(A^{r}\cap S^{\prime})=w_{H}(E_{t})\leq(1+\alpha)m(A^{l})\leq\frac{1+\alpha}{8}\mu(V),$
@@
where the last inequality follows from the assumption of the lemma. Putting everything together
$\mu(V\setminus S^{\prime})\geq\bar{m}(V\setminus S^{\prime})=\bar{m}(A^{r})-\bar{m}(A^{r}\cap S^{\prime})\geq\frac{\mu(V)}{2}-\frac{(1+\alpha)^{2}}{8}\mu(V)\geq\frac{1}{3}\mu(V)\geq\frac{1}{3}\mu(S^{\prime}).$
@@
$\Phi_{G}^{\mu}(S^{\prime},V\setminus S^{\prime})\mathrel{\mathop{:}}=\frac{|E_{G}(S^{\prime},V\setminus S^{\prime})|}{\min\{\mu(S^{\prime}),\mu(V\setminus S^{\prime})\}}\leq\frac{\frac{(1+\alpha)^{2}}{c}\mu(S^{\prime})}{\frac{1}{3}\mu(S^{\prime})}\leq\frac{7}{c}.$
As a direct corollary, we conclude that if Lemma 3.6 terminates in case $2$, then there exists a flow of constant congestion that routes one unit of flow from every edge in the cut $E(S,V\setminus S)$ into $S$, such that each vertex $v\in S$ receives at most $O(\phi\log n\cdot\mu(v))$ flow.
@@
Let $(S,V\setminus S)$ be the cut returned by Lemma 3.6, and assume $S\neq\emptyset$. Then there exists a flow of congestion $O(1)$ in $G[S]$ such that every $v\in S$ sends $\left(deg_{G}-deg_{G[S]}\right)(v)$ flow and receives at most $O(\phi\log n\cdot\mu(v))$ flow.
Recall that during the cut-matching game, the matching player produces cuts $S_{1},\ldots,S_{T}$ by applying Lemma 3.6. Since these cuts are disjoint, we get the following lemma.
Lemma A.4.
Let $(A,\bar{A})$ be the cut returned by case $2$ of Theorem 5.2. Then there exists a flow of congestion $O(1)$ in $G[\bar{A}]$ such that every $v\in\bar{A}$ sends $\left(deg_{G}-deg_{G[\bar{A}]}\right)(v)$ flow and receives at most $O(\phi\log n\cdot\mu(v))$ flow.
@@
Assume $|E[A,\bar{A})|>0$, as otherwise $A$ is a $(\phi,\mu)$-expander and we set $A^{\prime}=A$. Consider the following $(s,t)$-flow problem on a new graph $H=(V_{H}=A\cup\{s,t\},E_{H})$ constructed as follows. Start from $G$, and contract $V\setminus A$ into a single vertex, labeled as the source $s$ (ignoring self loops at $s$). Next, set the capacity of each edge ($G[A]$ and edges touching $s$) to be $\frac{3}{\phi}$. Finally, add a new sink vertex $t$ and connect it to each vertex $v\in A$ with an edge of capacity $\mu(v)$. Let $\alpha=0.1$, and compute a $(1+\alpha)$-fair cut $(S,T)$ of $H$ in time $\tilde{O}(m)$. Let $A^{\prime}=T\setminus\{t\}$. [^16] We later show that $A^{\prime}\neq\emptyset$. We now show that $A^{\prime}$ satisfies the requirements of the lemma.
First, suppose for contradiction that $G[A^{\prime}]$ is not a $\frac{\phi}{6}$-expander with respect to $\mu$. Then, there is a violating set $U\subseteq A^{\prime}$, satisfying $0<\mu(U)\leq\mu(A^{\prime}\setminus U)$ and
@@
Finally, we show the properties $\mu(A^{\prime})\geq\mu(A)-\frac{4}{\phi}|E(A,\bar{A})|$ and $|E(A^{\prime},\bar{A^{\prime}})|\leq 2|E(A,\bar{A})|$ promised by the lemma. Since $(S,T)$ is a $(1+\alpha)$-fair cut, it is in particular a $(1+\alpha)$-approximate $(s,t)$-mincut. Since $(\{s\},V_{H}\setminus\{s\})$ is an $(s,t)$-cut of capacity $\frac{3}{\phi}|E(A,\bar{A})|$, it follows that the cut $(S,T)$ has capacity at most $(1+\alpha)\frac{3}{\phi}\cdot|E(A,\bar{A})|$. To establish the first inequality of the theorem, observe that every vertex $v\in A\setminus A^{\prime}$ lies on the $S$-side of the cut $(S,T)$, and therefore contributes $\mu(v)$ to the capacity of $(S,T)$ through the edge $(v,t)$. Summing over all $v\in A\setminus A^{\prime}$, we obtain
$\mu(A\setminus A^{\prime})\leq|E(S,T)|\leq(1+\alpha)\cdot\frac{3}{\phi}|E(A,\bar{A})|\leq\frac{4}{\phi}|E(A,\bar{A})|~,$
which proves the first property. Moreover, since $\mu(A)\geq\frac{9|E(A,\bar{A})|\cdot}{\phi}$, it follows that $A^{\prime}\neq\emptyset$. For the second inequality, note that each edges $(u,v)$ in $E(A^{\prime},\bar{A^{\prime}})$ corresponds to an edge in $E(S,T)$ with capacity $\frac{3}{\phi}$, so summing over all such edges
$\frac{3}{\phi}|E(A^{\prime},\bar{A^{\prime}})|\leq|E(S,T)|\leq(1+\alpha)\frac{3}{\phi}|E(A,\bar{A})|~,$
meaning $|E(A^{\prime},\bar{A^{\prime}})|\leq(1+\alpha)|E(A,\bar{A})|$ which proves the second property.
@@
Let $(A^{\prime},\bar{A^{\prime}})$ be the cut returned from the trimming step (Theorem 5.1). The following flows exist.
@@
    A flow of congestion $\frac{3}{\phi}$ in $G[A^{\prime}]$, such that every $v\in A^{\prime}$ sends $\frac{2}{\phi}\left(deg_{G}-deg_{G[A^{\prime}]}\right)(v)$ flow and receives at most $\mu(v)$ flow. That is, every edge in the cut $E(A^{\prime},\bar{A^{\prime}})$ sends $\frac{2}{\phi}$ flow and every vertex $v\in A^{\prime}$ receives at most $\mu(v)$ flow.
@@
    A flow of congestion $O(1)$ in $G[\bar{A^{\prime}}]$ such that every $v\in\bar{A}$ sends $\left(deg_{G}-deg_{G[\bar{A}]}\right)(v)$ flow and receives at most $O(\phi\log n\cdot\mu(v))$ flow. That is, every edge in the cut $E(A^{\prime},\bar{A^{\prime}})$ sends one unit of flow and every vertex $v\in\bar{A}$ receives at most $O(\phi\log n\mu(v))$ flow. [^17] Note that vertices in $\bar{A^{\prime}}\setminus\bar{A}$ don’t receive any flow.

### v2

@@ Appendix A Omitted proofs
This appendix contains the proofs that were deferred from the main text. Section A.1 gives algebraic and embedding lemmas that support the analysis of the cut-matching game. In Section A.2 we introduce the fair-cut framework of [LNPSsoda13] and use it to analyze flow-based routines: first the flow subroutine in the cut-matching game (Lemma 3.6), and then the trimming step (Theorem 5.1).
@@ A.1 Algebraic and Embedding Proofs
Case $P_{L}\geq\frac{1}{20}P_{A_{t}}$: We set $\eta=0$, $A_{t}^{r}=R$, and $\bar{m}_{t}(i)=\mu(i)$ for $i\in A_{t}^{r}$. The construction of $A_{t}^{l}$ is as follows: If $\mu(L)\leq\mu(A_{t})/8$, then we set $A_{t}^{l}=L$, and $m_{t}(i)=\mu(i)$ for $i\in A_{t}^{l}$.
Otherwise, we sort $L$ according to $u_{i}$ and scan the vertices (from smallest to largest) until the accumulative $\mu(\cdot)$ weight of the scanned vertices is at least $\mu(A_{t})/8$. Let $i_{1},\ldots,i_{k}$ be the scanned vertices. We set $A_{t}^{l}=\{i_{1},\ldots,i_{k}\}$, $m_{t}(i_{j})=\mu(i_{j})$ for $1\leq i_{j}<k$, and $m_{t}(i_{k})=\mu(A_{t})/8-\mu(A_{t}^{l}\setminus\{i_{k}\})$. It is easy to verify that properties (1)(1)-$(3)$ of the lemma hold. Property $(4)$ holds because of the choice $\eta=0$. Finally, property $(5)$ holds since $P_{L}\geq\frac{1}{20}P_{A_{t}}$ and since $A_{t}^{l}$ contains the vertices with the smallest $u_{i}$ values (of total $\mu(\cdot)$ weight equaling $\mu(A_{t})/8$, while $\mu(L)\leq\mu(A_{t})/2$) in $L$.
@@
  • $A_{t}^{\ell}$ is the subset of $R^{\prime}$ consisting of the vertices with the largest $u_{i}$ values, chosen so that their total $\mu(\cdot)$ weight satisfies $\mu(A_{t}^{l})=\mu(A_{t})/8$. [^15] This is done similarly to the previous case. Again, $m_{t}(i)=\mu(i)$ for every $i\in A_{t}^{l}$, except for possibly one vertex.
@@ A.2 Fair Cuts
In this section, we prove Lemma 3.6 and Theorem 5.1, concerning the flow routines in the cut-matching step and the trimming step. The proofs are based on fair cuts, a stronger notion of approximate min-cut, introduced by [LNPSsoda13].
Definition A.1 (Fair Cut, [LNPSsoda13, li2025simple]).
@@
Theorem A.2 (Fair Cut, [LNPSsoda13, Theorem 1.1]).
@@
Let $\alpha=0.1$. Consider the following $(s,t)$-flow problem on a new graph $H=(V_{H}=V\cup\{s,t\},E_{H},w_{H})$ constructed as follows. We begin with $G$ and set its edge capacities to $\frac{c}{1+\alpha}$. We then add the source $s$ and connect it to every $v\in A^{l}$ with capacity $m(v)$. Finally, we add the sink $t$ and connect it to every $v\in A^{r}$ with capacity $\frac{1}{1+\alpha}\bar{m}(v)$.
@@
Using Theorem A.2, we compute (with high probability) a $(1+\alpha)$-fair cut $(S^{\prime},T^{\prime})$ with a corresponding feasible flow $f$ in $\tilde{O}(|E|)$ time. We split into the following two cases:
@@
Case 1: $S^{\prime}=\{s\}.$ By the definition of fair cut, for every $v\in A^{l}$, $f$ sends from $s$ to $v$ at least $\frac{1}{1+\alpha}m(v)$ units of flow. By decomposing $f$ into flow paths and reducing the flow sent on each path, we obtain a flow $f^{\prime}$ for $H$ that sends from $s$ to $v$ exactly $\frac{1}{1+\alpha}m(v)$ units of flow, for every $v\in A^{l}$. We claim that $f^{\prime\prime}=(1+\alpha)f^{\prime}$ is a feasible flow for $\Pi(G)$. Indeed, in $f^{\prime\prime}$, every $v\in A^{l}$ receives $m(v)$ units of flow from $s$, every $v\in A^{r}$ sends at most $(1+\alpha)\cdot\frac{1}{1+\alpha}\bar{m}(v)=\bar{m}(v)$ units of flow to $t$, and every $e\in E$ carries at most $(1+\alpha)\cdot\frac{1}{1+\alpha}c=c$ units of flow.
@@
Case 2: $S^{\prime}\neq\{s\}.$ We split the cut edges $E_{H}(S^{\prime},T^{\prime})$ into three sets: (i) the edges of $G$ crossing the cut: $E_{m}=E_{G}(S^{\prime}\setminus\{s\},T^{\prime}\setminus\{t\})$, (ii) the cut edges incident to $s$: $E_{s}=\{(s,v)\in E_{H}\mid v\in T^{\prime}\}$, and (iii) the cut edges incident to $t$: $E_{t}=\{(v,t)\in E_{H}\mid v\in S^{\prime}\}$. Let $\bar{E}_{s}$ be the incident edges to $s$ that are not in $E_{s}$.
@@
We decompose $f$ into paths. By the definition of a fair cut, every edge $(s,v)\in E_{s}$ sends at least $\frac{1}{1+\alpha}m(v)$ units of flow. We scale down the flow in the flow paths as follows. We delete all flow paths that start with an edge of $\bar{E}_{s}$. The rest of the flow paths are scaled down until each edge $(s,v)\in E_{s}$ sends exactly $\frac{1}{1+\alpha}m(v)$ units of flow. This results in a flow $f^{\prime}$. The algorithm then returns $f^{\prime\prime}=(1+\alpha)f^{\prime}$ and the cut $S=S^{\prime}\setminus\{s\}$.
@@
Similarly to the previous case, $f^{\prime\prime}$ is a feasible flow for the flow problem $\Pi(G\setminus S)$. It remains to prove that $\Phi_{G}^{\mu}(S,V\setminus S)\mathrel{\mathop{:}}=\frac{|E_{m}|}{\min\{\mu(S),\mu(V\setminus S)\}}\leq\frac{7}{c}$. We begin by upper bounding the numerator. For the remainder of the analysis, we consider again the flow $f$ and its path decomposition. By the definition of a fair cut, every flow path that traverses through $E_{m}$ must start with an edge in $\bar{E}_{s}$, and every edge in $E_{m}$ is at least $\frac{1}{1+\alpha}$-saturated. Therefore,
$\frac{c}{1+\alpha}\cdot|E_{G}(S,V\setminus S)|=w_{H}(E_{m})\leq(1+\alpha)w_{H}(\bar{E}_{s})=(1+\alpha)m_{t}(S)\leq(1+\alpha)\mu(S).$ (5)
@@
To bound the denominator we prove that $\mu(V\setminus S)=\Omega(\mu(S))$. Recall the assumption of the lemma $\bar{m}(A^{r})\geq\mu(V)/2$. We prove that $\bar{m}(A^{r}\cap S)$ is small. Indeed, for every $v\in A^{r}\cap S$, $(v,t)\in E_{t}$ is crossing the cut $(S^{\prime},T^{\prime})$ and therefore is $\frac{1}{1+\alpha}$-saturated. Since the total flow crossing $(S^{\prime},T^{\prime})$ is bounded by the total source capacity (which is $m(A^{l})$), we get that
$\frac{1}{1+\alpha}\bar{m}(A^{r}\cap S)=w_{H}(E_{t})\leq(1+\alpha)m(A^{l})\leq\frac{1+\alpha}{8}\mu(V),$
@@
where the last inequality follows from the assumption of the lemma. Putting everything together,
$\mu(V\setminus S)\geq\bar{m}(V\setminus S)=\bar{m}(A^{r})-\bar{m}(A^{r}\cap S)\geq\frac{\mu(V)}{2}-\frac{(1+\alpha)^{2}}{8}\mu(V)\geq\frac{1}{3}\mu(V)\geq\frac{1}{3}\mu(S).$
@@
$\Phi_{G}^{\mu}(S,V\setminus S)\mathrel{\mathop{:}}=\frac{|E_{G}(S,V\setminus S)|}{\min\{\mu(S),\mu(V\setminus S)\}}\leq\frac{\frac{(1+\alpha)^{2}}{c}\mu(S)}{\frac{1}{3}\mu(S)}\leq\frac{7}{c}.$
As a direct corollary, we observe that if Lemma 3.6 terminates with Case (2), we can “break” the flow paths at the (fair-)cut $(S,V\setminus S)$ and obtain two flows satisfying the following guarantees.
@@
Let $(S,V\setminus S)$ be the cut returned by Lemma 3.6, and assume $S\neq\emptyset$. Then:
  • There exists a flow of congestion $O(1)$ in $G[S]$ such that every $v\in S$ sends $\left(\deg_{G}-\deg_{G[S]}\right)\!(v)$ units of flow and receives at most $O(\phi\log n\cdot\mu(v))$ units of flow.
  • There exists a flow of congestion $O(1)$ in $G[V\setminus S]$ such that every $v\in V\setminus S$ sends $\left(\deg_{G}-\deg_{G[V\setminus S]}\right)\!(v)$ units of flow and receives at most $O(\phi\log n\cdot\mu(v))$ units of flow.
Proof.
In the notation of the proof of Lemma 3.6, the assumption $S\neq\emptyset$ corresponds to Case (2) of the proof. Let $f$ be the flow returned by the fair-cut routine, and recall $E_{m}=E_{G}(S^{\prime}\setminus\{s\},T^{\prime}\setminus\{t\})=E_{G}(S,V\setminus S),$ the set of edges of $G$ that cross the fair-cut. Recall that $f$ is routed in $H$, the auxiliary graph, with constant congestion. In particular, $f$ may be routed in $G$ with congestion $O(c)$.
To construct the two flows required by the corollary, we proceed as follows. Decompose $f$ into flow paths and remove the paths that do not traverse any edge of $E_{m}$. Denote the remaining flow by $f_{1}$. By the definition of a fair-cut, every edge $(u,v)\in E_{m}$ is at least $\frac{1}{1+\alpha}$-saturated (recall $\alpha=0.1$). Hence $f_{1}$ routes at least $\frac{1}{1+\alpha}c=\Theta\left(\frac{1}{(1+\alpha)\phi\log n}\right)$ units of flow through each $(u,v)\in E_{m}$ in the direction from $S$ to $V\setminus S$. We then scale down the flow paths of $f_{1}$ so exactly $\frac{1}{1+\alpha}c$ units of flow are routed on each edge of $E_{m}$. Denote this scaled flow by $f_{2}$. Finally, define $f_{3}\mathrel{\mathop{:}}=\frac{1+\alpha}{c}f_{2}$. Then $f_{3}$ routes exactly one unit of flow across each edge in $E_{m}$ and has congestion $O(1)$ in $G$. Moreover, for every $v\in V$ the source and sink capacities at $v$ in $f_{3}$ are bounded by $O\left(\frac{1+\alpha}{c}\mu(v)\right)=O(\phi\log n\mu(v))$.
Observe that $f_{3}$ consists of flow paths that cross each edge of the cut $(S,V\setminus S)$ once. We obtain the two desired flows as follows:
  • The first flow is obtained by taking the prefix of each path until the crossing edge, deleting $s$ and then reversing the path so that the flow is routed from the cut $E_{m}=E_{G}(S,V\setminus S)$ to $S$.
  • The second flow is obtained by taking the suffix of each path that starts at the crossing edge, deleting $t$.
Recall that during the cut-matching game, the matching player produces cuts $S_{1},\ldots,S_{t}$ (up to some $t<T$) by applying Lemma 3.6. Since these cuts are disjoint, we get the following corollary.
Corollary A.4.
Let $(A,\overline{A})$ be the cut returned by Case (2) or Case (3) of Theorem 3.1. Then, there exists a flow of congestion $O(1)$ in $G[\overline{A}]$ such that every $v\in\overline{A}$ sends $\left(\deg_{G}-\deg_{G[\overline{A}]}\right)\!(v)$ units of flow and receives at most $O(\phi\log n\cdot\mu(v))$ units of flow.
@@
Assume $|E(A,\overline{A})|>0$, as otherwise $A$ is a $(\phi,\mu)$-expander and we set $A^{\prime}=A$. Consider the following $(s,t)$-flow problem on a new graph $H=(V_{H}=A\cup\{s,t\},E_{H})$ constructed as follows. Start from $G$, and contract $V\setminus A$ into a single vertex, labeled as the source $s$ (ignoring self loops at $s$). Next, set the capacity of each edge ($G[A]$ and edges touching $s$) to be $\frac{3}{\phi}$. Finally, add a new sink vertex $t$ and connect it to each vertex $v\in A$ with an edge of capacity $\mu(v)$. Let $\alpha=0.1$, and compute a $(1+\alpha)$-fair cut $(S,T)$ of $H$ in time $\tilde{O}(m)$. Let $A^{\prime}=T\setminus\{t\}$. [^17] We later show that $A^{\prime}\neq\emptyset$. We now show that $A^{\prime}$ satisfies the requirements of the lemma.
First, suppose for contradiction that $G[A^{\prime}]$ is not a $\frac{\phi}{6}$-expander with respect to $\mu$. Then, there is a violating set $U\subseteq A^{\prime}$ (see Figure 1), satisfying $0<\mu(U)\leq\mu(A^{\prime}\setminus U)$ and
@@
Finally, we show the properties $\mu(A^{\prime})\geq\mu(A)-\frac{4}{\phi}|E(A,\overline{A})|$ and $|E(A^{\prime},\overline{A^{\prime}})|\leq 2|E(A,\overline{A})|$ promised by the lemma. Since $(S,T)$ is a $(1+\alpha)$-fair cut, it is in particular a $(1+\alpha)$-approximate $(s,t)$-mincut. Since $(\{s\},V_{H}\setminus\{s\})$ is an $(s,t)$-cut of capacity $\frac{3}{\phi}|E(A,\overline{A})|$, it follows that the cut $(S,T)$ has capacity at most $(1+\alpha)\frac{3}{\phi}\cdot|E(A,\overline{A})|$. To establish the first inequality of the theorem, observe that every vertex $v\in A\setminus A^{\prime}$ lies on the $S$-side of the cut $(S,T)$, and therefore contributes $\mu(v)$ to the capacity of $(S,T)$ through the edge $(v,t)$. Summing over all $v\in A\setminus A^{\prime}$, we obtain
$\mu(A\setminus A^{\prime})\leq|E(S,T)|\leq(1+\alpha)\cdot\frac{3}{\phi}|E(A,\overline{A})|\leq\frac{4}{\phi}|E(A,\overline{A})|~,$
which proves the first property. Moreover, since $\mu(A)\geq\frac{9|E(A,\overline{A})|\cdot}{\phi}$, it follows that $A^{\prime}\neq\emptyset$. For the second inequality, note that each edges $(u,v)$ in $E(A^{\prime},\overline{A^{\prime}})$ corresponds to an edge in $E(S,T)$ with capacity $\frac{3}{\phi}$, so summing over all such edges
$\frac{3}{\phi}|E(A^{\prime},\overline{A^{\prime}})|\leq|E(S,T)|\leq(1+\alpha)\frac{3}{\phi}|E(A,\overline{A})|~,$
meaning $|E(A^{\prime},\overline{A^{\prime}})|\leq(1+\alpha)|E(A,\overline{A})|$ which proves the second property.
@@
Let $(A^{\prime},\overline{A^{\prime}})$ be the cut returned from the trimming step (Theorem 5.1) in the proof of Theorem 5.2. Then, the following flows exist.
@@
    A flow of congestion $O(1)$ in $G[A^{\prime}]$, such that every $v\in A^{\prime}$ sends $\left(\deg_{G}-\deg_{G[A^{\prime}]}\right)\!(v)$ units of flow and receives at most $O(\phi\cdot\mu(v))$ units of flow.
@@
    A flow of congestion $O(1)$ in $G[\overline{A^{\prime}}]$ such that every $v\in\overline{A^{\prime}}$ sends $\left(\deg_{G}-\deg_{G[\overline{A^{\prime}}]}\right)\!(v)$ units of flow and every $v\in\overline{A}$ receives at most $O(\phi\log n\cdot\mu(v))$ units of flow. [^18] Note that vertices in $\overline{A^{\prime}}\setminus\overline{A}$ don’t receive any flow.